# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

### v3

@@ Abstract
Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past years. It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques. Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT. MToP incorporates over 40 MTEAs, more than 150 MTO problem cases with real-world applications, and over 20 performance metrics. Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 40 popular single-task evolutionary algorithms to address MTO problems. MToP boasts a user-friendly graphical interface, facilitating results analysis, data export, and schematics visualization. More importantly, MToP is designed with extensibility in mind, allowing users to develop new algorithms and tackle emerging problem domains. The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
Index Terms:
@@
Evolutionary multitasking, MATLAB optimization platform, multitask optimization problem, evolutionary algorithm.
I Introduction
@@
Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency. Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and user-friendly nature. Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization [1, 2], multi-objective optimization [3, 4], and combinatorial optimization [5, 6]. In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking [7]. A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ optimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
$\begin{gathered}(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})=\arg\min\Big{[}F_{1}(\bm{x}_{1}),F_{2}(\bm{x}_{2}),...,F_{K}(\bm{x}_{K})\Big{]},\\ \textit{s.t. }\bm{x}_{k}\in\Omega_{k},\ k=1,2,...,K,\\ \end{gathered}$ (1)
@@
where $\Omega_{k}$ and $F_{k}$ are the decision space and objective function of the $k$-th task. Note that for a multi-objective optimization task, the $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task. For task $k$ with upper and lower bounds $L_{k}$ and $U_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space [8] as follows:
$\bm{x}^{\prime}=\frac{\bm{x}-L_{k}}{U_{k}-L_{k}}.$ (2)
@@
In addition, the dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$. During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision. EMT has found successful applications in various domains, including engineering scheduling [5, 9], nonlinear equation systems [10], feature selection in machine learning [11], anomaly detection [12], point cloud registration [13, 14], and reinforcement learning [15].
@@
To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques [16, 17]. Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources [18, 19]. The first attempt of EMT can be traced back to the multifactorial EA [8], which introduced an implicit knowledge representation via random mating among optimization tasks. Subsequently, numerous EAs tailored for MTO have emerged. These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks [20, 21], or a multi-population framework allocating separate populations for each task [22, 23]. Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space [8, 21], autoencoding [24, 25], affine transformation [26, 27, 28], and adversarial generative models [29] have been proposed in MTEAs. Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment [20, 30, 31], knowledge selection [23, 32], and historical feedback [33, 34, 35], have also been investigated. Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes. This presents challenges for newcomers entering the field of EMT and for researchers seeking to conduct convenient experimental comparisons of algorithms.
@@
As the field of EMT continues to gain momentum, there is an urgent need for a convenient and user-friendly software platform to facilitate the benchmarking of MTEAs. Furthermore, accessible source code and platforms are indispensable for exploring the real-world applications of the EMT field. Open-source and user-friendly software platforms play a crucial role in fostering the advancement of a research field. In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization. For instance, PlatEMO [36] is tailored for evolutionary multi-objective optimization and benchmarking multi-objective EAs, aiding researchers in designing EAs to address various types of multi-objective optimization problems. EvoX [37] is a distributed GPU-accelerated library that assists researchers in designing parallel EAs to expedite the resolution of complex optimization problems and reinforcement learning training. Evo-Panel [38] created a comprehensive dynamic visualization tool to illustrate the detailed procedures of different optimization algorithms in solving numerical benchmark functions. EDOLAB [39] is an optimization tool specifically for dynamic optimization.
@@
However, unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks. Especially important, the implementation of MTEAs requires incorporating inter-task solution space mapping and diverse knowledge transfer techniques. Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics. Specifically, single-task metrics evaluate the performance of each task in the MTO problem, while multitask metrics assess the overall performance of the MTO problem. These bring uncertainty to the pattern of results display and analysis. These distinctive requirements present significant challenges for the implementation of MTEAs in existing tools and platforms. Moreover, EMT has been extended to the subfields of many-task optimization [40, 41], multi-objective multitask optimization [42, 43], competitive multitask optimization [33, 44], and constrained multi-task optimization [45]. Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
@@
  2. 2.
    MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API. Over 40 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types. Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 40 single-task EAs of diverse types. In terms of synthetic test problems, MToP incorporates over 150 benchmark MTO problems alongside several real-world applications. Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities.
  3. 3.
    MToP is designed to be easily extended, allowing for the seamless addition of new algorithms, problems, and metrics. By adhering to established coding patterns and implementing functionality based on the public API, new code can be seamlessly integrated and utilized within MToP. Given its status as a completely open-source project, researchers have the opportunity to leverage existing algorithms, problems, and metrics as the foundation for novel ideas. Through the collaborative platform GitHub, MToP undergoes continuous updates and enhancements, ensuring that it remains at the forefront of EMT research [^1] MToP has been continuously updated since Sep. 21, 2021, when it was uploaded to GitHub at https://github.com/intLyc/MTO-Platform..
The rest of this paper is organized as follows. Section II elaborates on the architecture of MToP. Section III describes approaches for performing experiments with MToP. Section IV introduces how to extend the contents of MToP. Finally, the conclusion and future work are given in Section V.
II Architecture of MToP
@@
II-A Functional Modules
II-A1 Test Module
@@
Figure 1: Examples of graphical display in the test module of MToP. (a) and (b) illustrate the landscapes of single-objective problems with different tasks in the one- and two-dimensional unified search space, respectively. (c) shows the feasible regions of a single-objective problem with different tasks in the two-dimensional unified search space. (d) depicts the Pareto front of a multi-objective problem with multiple tasks. (e) displays the convergence behavior of metrics after executing algorithms on problems.
@@
Figure 2: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space and decision space with the number of generations. The top row (a)-(f) shows the population in the objective space, while the bottom row (g)-(l) displays the population in the decision space.
@@
Figure 3: The application programming interface calling workflow of running an algorithm MFEA on a problem CEC17-MTP1 with metric Obj by MToP.
@@
II-A2 Experiment Module
TABLE I: Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
@@ II-A2 Experiment Module
Figure 4: schematic plotting by the experiment module of MToP. (a) Convergence plot of algorithms on problems. (b) Pareto front plot of multi-objective problems.
@@
Metrics functions operating on the data are categorized into single-task and multitask metrics. The selection of single-task metrics presents the results of each task of the MTO problem in tabular format within MToP while choosing to multitask metrics displays the results of each MTO problem. Significance test results are also included in the table alongside the metric results. All table data can be exported in formats such as tex, csv, and xlsx, as demonstrated in Table I.
@@
II-A3 Data Process Module
TABLE II: Properties of experimental data.
@@ II-A3 Data Process Module
Each unique execution within the experiment module is considered a data object and is saved to a data file named <MTOData.mat>. The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table II. This saved data can subsequently be reloaded by the experiment module of MToP.
@@
II-B File Structure
Figure 5: File structure of MToP. The root directory contains the main script file <mto.m> and four subfolders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
@@ II-B File Structure
II-C Code Pattern
II-C1 Algorithm
TABLE III: Properties and methods of algorithm base class.
@@ II-C1 Algorithm
All algorithms within MToP inherit from the algorithm base class <Algorithm.m>. The properties and methods of this base class are detailed in Table III. This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself. The MTEAs implemented in MToP are outlined in Table LABEL:supp:tab:algo-mt of the supplementary files. Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems. Specific algorithms belonging to this category are listed in Table LABEL:supp:tab:algo-st.

### v4

@@ Abstract.
Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past decade. It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques. Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT. MToP incorporates over 50 MTEAs, more than 200 MTO problem cases with real-world applications, and over 20 performance metrics. Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 50 popular single-task evolutionary algorithms to address MTO problems. MToP boasts a user-friendly graphical interface, facilitating results analysis, data export, and schematics visualization. More importantly, MToP is designed with extensibility in mind, allowing users to develop new algorithms and tackle emerging problem domains. The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
@@
Evolutionary multitasking, MATLAB optimization platform, multitask optimization problem, evolutionary algorithm
1. Introduction
@@
Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency. Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and user-friendly nature. Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization (Wang et al., 2020; Tian et al., 2021b), multi-objective optimization (Zhang and Li, 2007; Deb and Jain, 2014), and combinatorial optimization (Feng et al., 2021c, a). In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking (Wei et al., 2022). A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ optimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
(1) $\begin{gathered}(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})=\arg\min\Big{[}F_{1}(\bm{x}_{1}),F_{2}(\bm{x}_{2}),...,\bm{F}_{k}(\bm{x}_{K})\Big{]},\\ \textit{s.t. }\bm{x}_{k}\in\Omega_{k},\ k=1,2,...,K,\\ \end{gathered}$
@@
where $\Omega_{k}$ and $\bm{F}_{k}$ are the decision space and objective function of the $k$-th task. Note that for a multi-objective optimization task, the $\bm{F}_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task. For task $k$ with upper and lower bounds $\bm{L}_{k}$ and $\bm{U}_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space (Gupta et al., 2016) as follows:
(2) $\bm{x}^{\prime}=\frac{\bm{x}-\bm{L}_{k}}{\bm{U}_{k}-\bm{L}_{k}}.$
@@
In addition, the dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$. During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision. EMT has found successful applications in various domains, including engineering scheduling (Feng et al., 2021c; Gupta et al., 2022), nonlinear equation systems (Li et al., 2024e), feature selection in machine learning (Chen et al., 2022), anomaly detection (Wang et al., 2022b), point cloud registration (Wu et al., 2024a, b), and reinforcement learning (Zhang et al., 2023).
@@
To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques (Tan et al., 2023; Gupta et al., 2018). Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources (Tan et al., 2021; Li et al., 2024f). The first attempt of EMT can be traced back to the multifactorial EA (Gupta et al., 2016), which introduced an implicit knowledge representation via random mating among optimization tasks. Subsequently, numerous EAs tailored for MTO have emerged. These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks (Bali et al., 2020; Zhou et al., 2021b), or a multi-population framework allocating separate populations for each task (Li et al., 2020; Wang et al., 2022a). Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space (Gupta et al., 2016; Zhou et al., 2021b), autoencoding (Feng et al., 2019b; Zhou et al., 2021a), affine transformation (Xue et al., 2022; Lin et al., 2024; Gao et al., 2022), and adversarial generative models (Liang et al., 2023) have been proposed in MTEAs. Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment (Bali et al., 2020; Jiang et al., 2023a; Wu et al., 2024c), knowledge selection (Wang et al., 2022a; Liang et al., 2022c), and historical feedback (Li et al., 2022c; Lin et al., 2021; Tang et al., 2021b), have also been investigated. Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes. This presents challenges for newcomers entering the field of EMT and for researchers seeking to conduct convenient experimental comparisons of algorithms.
@@
As the field of EMT continues to gain momentum, there is an urgent need for a convenient and user-friendly software platform to facilitate the benchmarking of MTEAs. Furthermore, accessible source code and platforms are indispensable for exploring the real-world applications of the EMT field. Open-source and user-friendly software platforms play a crucial role in fostering the advancement of a research field. In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization. For instance, PlatEMO (Tian et al., 2017) is tailored for evolutionary multi-objective optimization and benchmarking multi-objective EAs, aiding researchers in designing EAs to address various types of multi-objective optimization problems. EvoX (Huang et al., 2024) is a distributed GPU-accelerated library that assists researchers in designing parallel EAs to expedite the resolution of complex optimization problems and reinforcement learning training. EDOLAB (Peng et al., 2023) is an optimization tool specifically for dynamic optimization.
@@
However, unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks. Especially important, the implementation of MTEAs requires incorporating inter-task solution space mapping and diverse knowledge transfer techniques. Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics. Specifically, single-task metrics evaluate the performance of each task in the MTO problem, while multitask metrics assess the overall performance of the MTO problem. These bring uncertainty to the pattern of results display and analysis. These distinctive requirements present significant challenges for the implementation of MTEAs in existing tools and platforms. Moreover, EMT has been extended to the subfields of many-task optimization (Chen et al., 2020; Liaw and Ting, 2019), multi-objective MTO (Gupta et al., 2017; Gao et al., 2023), competitive MTO (Li et al., 2022c, 2023a), and constrained multi-task optimization (Li et al., 2022a). Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
@@
  2. (2)
    MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API. Over 50 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types. Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 50 single-task EAs of diverse types. In terms of synthetic test problems, MToP incorporates over 200 benchmark MTO problems alongside several real-world applications. Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities.
  3. (3)
    MToP is designed to be easily extended, allowing for the seamless addition of new algorithms, problems, and metrics. By adhering to established coding patterns and implementing functionality based on the public API, new code can be seamlessly integrated and utilized within MToP. Given its status as a completely open-source project, researchers have the opportunity to leverage existing algorithms, problems, and metrics as the foundation for novel ideas. Through the collaborative platform GitHub, MToP undergoes continuous updates and enhancements, ensuring that it remains at the forefront of EMT research.
The rest of this paper is organized as follows. Section 2 elaborates on the architecture of MToP. Section 3 describes approaches for performing experiments with MToP. Section 4 introduces how to extend the contents of MToP. Finally, the conclusion and future work are given in Section 5.
2. Architecture of MToP
@@
2.1. Functional Modules
2.1.1. Test Module
@@
Figure 1. Examples of graphical display in the test module of MToP. (a) and (b) illustrate the landscapes of single-objective problems with different tasks in the one- and two-dimensional unified search space, respectively. (c) shows the feasible regions of a single-objective problem with different tasks in the two-dimensional unified search space. (d) depicts the Pareto front of a multi-objective problem with multiple tasks. (e) displays the convergence behavior of metrics after executing algorithms on problems.
@@
Figure 2. Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space and decision space with the number of generations. The top row (a)-(f) shows the population in the objective space, while the bottom row (g)-(l) displays the population in the decision space.
@@
Figure 3. The application programming interface calling workflow of running an algorithm MFEA on a problem CEC17-MTP1 with metric Obj by MToP.
@@
2.1.2. Experiment Module
Table 1. Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
@@ 2.1.2. Experiment Module
Figure 4. schematic plotting by the experiment module of MToP. (a) Convergence plot of algorithms on problems. (b) Pareto front plot of multi-objective problems.
@@
Metrics functions operating on the data are categorized into single-task and multitask metrics. The selection of single-task metrics presents the results of each task of the MTO problem in tabular format within MToP while choosing to multitask metrics displays the results of each MTO problem. Significance test results are also included in the table alongside the metric results. All table data can be exported in formats such as tex, csv, and xlsx, as demonstrated in Table 1.
@@
2.1.3. Data Process Module
Table 2. Properties of experimental data.
@@ 2.1.3. Data Process Module
Each unique execution within the experiment module is considered a data object and is saved to a data file named <MTOData.mat>. The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table 2. This saved data can subsequently be reloaded by the experiment module of MToP.
@@
2.2. File Structure
Figure 5. File structure of MToP. The root directory contains the main script file <mto.m> and four subfolders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
@@ 2.2. File Structure
2.3. Code Pattern
2.3.1. Algorithm
Table 3. Properties and methods of algorithm base class.
@@ 2.3.1. Algorithm
All algorithms within MToP inherit from the algorithm base class <Algorithm.m>. The properties and methods of this base class are detailed in Table 3. This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself. The MTEAs implemented in MToP are outlined in Table 4. Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems. Specific algorithms belonging to this category are listed in Table 5.